# Universal Approximation Theory: The basic theory for deep learning-based computer vision models

### v2

@@ A.1 The Properties of UAT
Since the first term of $\mathbf{x}^{\prime}_{i+1}$ has the same mathematical form as UAT, and the second term can be considered as a component within UAT, the sum of both terms still corresponds to the UAT mathematical form. Therefore, we have proven that the mathematical form of a multi-layer residual network is consistent with the UAT mathematical form. The difference lies in that some bias terms are approximated using UAT rather than being directly defined.
A.2 The UAT Format of Residual-based CNNs
To demonstrate its relationship with UAT, we will provide its general form based on the definitions in Figure 8 in main text and then compare it with the general form presented in Section A.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$. By examining the relationship between these two forms, we can prove that the mathematical form of a multi-layer residual-based convolutional network is consistent with the UAT. The general term of residual-based CNNs is shown in Eq. 7
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{x}_{i}^{\prime}+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}^{\prime}_{i}+\mathbf{b}_{i+1,1})+\mathbf{b}_{i+1,2}$ (7)
$=(\mathbf{x}_{i}^{\prime}+\mathbf{b}_{i+1,2})+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}^{\prime}_{i}+\mathbf{b}_{i+1,1})$
Compared to $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, the only difference lies in $\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3}$. It is easy to deduce that the presence or absence of $\mathbf{W}^{\prime}_{i+1,1}$ does not affect the overall mathematical form. Therefore, residual-based CNNs are also concrete implementations of the UAT. To clearly present the UAT format, we will use the method mentioned in Section A.1. Assuming the mathematical form of $\mathbf{x}^{\prime}_{i}$ is consistent with UAT, we decompose $\mathbf{x}^{\prime}_{i}$ into a primary term plus a residual term, written as $\mathbf{x}^{\prime}_{i}=(\mathbf{x}_{0}+\mathbf{b}_{i,2})+UAT^{R}_{i}$. Substituting this into $(\mathbf{x}_{i}^{\prime}+\mathbf{b}_{i+1,2})+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}^{\prime}_{i}+\mathbf{b}_{i+1,1})$, we obtain Eq. (9). Setting $\mathbf{b}_{i+1,2}=\underline{\mathbf{b}_{i,2}+\mathbf{b}_{i+1,2}}$ and $\mathbf{b}_{i+1,1}=\underline{(\mathbf{W}_{i+1,1}^{\prime}\mathbf{b}_{i,1}+\mathbf{b}_{i+1,1})+\mathbf{W}_{i+1,1}^{\prime}UAT^{R}_{i}}$, we finally get $\mathbf{x}_{i+1}=(\mathbf{x}_{0}+\mathbf{b}_{i+1,2})+UAT^{R}_{i}+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})$. Since the mathematical form of $UAT^{R}_{i}$ is consistent with multi-layer UAT, it follows that multi-layer residual CNNs are also specific implementations of UAT. In addition, $UAT^{R}_{i+1}=UAT^{R}_{i}+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})$. So, $\mathbf{x}_{i+1}=(\mathbf{x}_{0}+\mathbf{b}_{i+1,2})+UAT^{R}_{i+1}$, where $UAT_{i+1}^{R}=\Sigma_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,2}\sigma(\mathbf{W}^{\prime}_{j,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,1})$, $\mathbf{b}^{\prime}_{j,1}=(\mathbf{W}^{\prime}_{j,1}\mathbf{b}^{\prime}_{j-1,2}+\mathbf{b}^{\prime}_{j,1})+\mathbf{W}^{\prime}_{j,1}UAT_{j-1}^{R}$. In conclusion, the UAT format of $i+1$-layer residual-based CNN is:
@@ A.2 The UAT Format of Residual-based CNNs
$\mathbf{b}^{\prime}_{j,1}$ is a $j$-layer UAT and the input is $\mathbf{x}_{0}$. Of course, aside from the parameter $\mathbf{b}^{\prime}_{j,1}$, which changes as the number of layers increases, other parameters will also vary with the number of layers. However, as demonstrated in the above derivation, these changes involve only matrix multiplications or additions. Thus, they do not fundamentally alter the mathematical form of the UAT. Therefore, we will not detail the specific forms of them here. But we provide some example in Figure 1
@@
Figure 1: Some examples of the UAT format of multi-layer residual-based convolution.)
A.3 The UAT Format of Transformer-based ViTs
Similarly, To illustrate its relationship with the UAT, we will first derive the general form based on Figure 9 in main text and then compare it with the general form given in Section A.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2}).$
By examining the relationship between these forms, we can demonstrate that the mathematical form of multi-layer Transformer networks is consistent with the UAT. The Transformer involves two key operations: multi-head attention (MHA) and feed-forward networks (FFN). These operations, when expressed in matrix-vector form, correspond to Eq.(10) and Eq.(11), respectively. Consequently, the general term for a Transformer-based network can be written as Eq. (12).
$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (10)
$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (11)
@@ A.3 The UAT Format of Transformer-based ViTs
Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}$, so we have $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$. It is evident that, when compared to the general form given in Section A.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, the mathematical forms are the same. Therefore, multi-layer Transformers are also specific implementations of the UAT. Similarly, to clearly present the UAT format, we use the method mentioned in Section A.1. Assuming the mathematical form of $\mathbf{x}^{\prime}_{i}$ is consistent with UAT, we decompose $\mathbf{x}^{\prime}_{i}$ into a primary term plus a residual term, written as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,3})+UAT^{R}_{i}$. Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we obtain Eq. (14). By setting $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,3}=\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}$, $UAT^{R}_{i}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$, and $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$, we get $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,3})+UAT^{R}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,2})$. Since the mathematical form of $UAT^{R}_{i}$ is consistent with multi-layer UAT, it follows that multi-layer Transformer networks are also specific implementations of UAT. In addition, $UAT^{R}_{i+1}=UAT^{R}_{i}+\mathbf{W}_{i+1,3}^{\prime}\sigma(\mathbf{W}_{i+1,2}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,2})$. So, $\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3})+UAT^{R}_{i+1}$, where $i=0,1,2\ldots$, $UAT_{i+1}^{R}=\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$. To sum up, the $i+1$-layer Transformer can be written as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (13)
The same to residual-based CNNs, other parameters will also vary with the number of layers. And, these changes involve only matrix multiplications or additions. Thus, they do not fundamentally alter the mathematical form of the UAT. Therefore, we will not detail the specific forms for each layer here.
The biggest difference between Transformers and CNNs lies in the fact that the parameters in the Multi-Head Attention (MHA) mechanism dynamically change with the input. Therefore, in the corresponding UAT mathematical form for Transformers, all $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ parameters for the $i$-th layer, where $j=1,...,i$, are dynamically changing with the input. We give some examples in Figure 2.
@@
Figure 2: Some examples of the UAT format of multi-layer residual-based convolution.)

### v4

@@ A.2 From UAT to DUAT
Let $UAT^{R}_{i+1}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$. Then, we can express $\mathbf{x}^{\prime}_{i+1}$ as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,1})+UAT^{R}_{i+1}$. Thus, apart from the initial term, the overall mathematical form of $\mathbf{x}^{\prime}_{i+1}$ remains consistent with UAT. Since a single term does not alter the nature of UAT, we conclude that the overall form of a multi-layer residual network aligns with UAT. Furthermore, as certain bias parameters within the network, such as $\mathbf{b}^{\prime}_{i+1,2}$, are also approximated using UAT, we have therefore demonstrated that the mathematical form of a multi-layer residual network is a UAT function.
A.3 The DUAT Format of Residual-based CNNs
To demonstrate that the residual-based CNN is indeed a DUAT function, we present its general form as defined by Figure 8 in the main text, represented by Eq. 16. Comparing this with the general form provided in Section A.2, specifically $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we find their only distinction lies in the term $\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3}$. It is straightforward to deduce that the presence or absence of $\mathbf{W}^{\prime}_{i+1,1}$ does not affect the overall mathematical form. Therefore, the residual-based CNN also qualifies as a DUAT function.
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{x}_{i}^{\prime}+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}^{\prime}_{i}+\mathbf{b}_{i+1,1})+\mathbf{b}_{i+1,2}^{\prime}$ (16)
$=(\mathbf{x}_{i}^{\prime}+\mathbf{b}_{i+1,2}^{\prime})+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}^{\prime}_{i}+\mathbf{b}_{i+1,1}^{\prime})$
To clearly present the DUAT form of residual-based CNNs, we apply the method described in Section A.2. Assuming that the general mathematical form of $\mathbf{x}^{\prime}_{i}$ aligns with UAT, we decompose $\mathbf{x}^{\prime}_{i}$ into a primary term and a remainder term, expressed as $\mathbf{x}^{\prime}_{i}=(\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i,2}^{\prime})+UAT^{R}_{i}$, where $UAT_{i}^{R}=\sum_{j=1}^{i}\mathbf{W}^{\prime}_{j,2}\sigma(\mathbf{W}^{\prime}_{j,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,1})$ for $i\geq 1$. Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{x}_{i}^{\prime}+\mathbf{b}_{i+1,2}^{\prime})+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}^{\prime}_{i}+\mathbf{b}_{i+1,1}^{\prime})$, we derive Eq. 18.
By setting $\mathbf{b}_{i+1,2}^{\prime}=\mathbf{b}_{i,2}^{\prime}+\mathbf{b}_{i+1,2}^{\prime}$ and $\mathbf{b}_{i+1,1}^{\prime}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{b}_{i,1}^{\prime}+\mathbf{b}_{i+1,1}^{\prime})+\mathbf{W}_{i+1,1}^{\prime}UAT^{R}_{i}$, we obtain $\mathbf{x}_{i+1}^{\prime}=(\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,2}^{\prime})+UAT^{R}_{i}+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,1}^{\prime})$. Given that the overall mathematical format of $UAT^{R}_{i}$ is consistent with the UAT formula, it follows that $\mathbf{x}_{i+1}^{\prime}$ also conforms to the UAT format.
Furthermore, we define $UAT^{R}_{i+1}=UAT^{R}_{i}+\mathbf{W}_{i+1,2}^{\prime}\sigma(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,1}^{\prime})$, leading to $\mathbf{x}_{i+1}^{\prime}=(\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,2}^{\prime})+UAT^{R}_{i+1}$, where $UAT_{i+1}^{R}=\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,2}\sigma(\mathbf{W}^{\prime}_{j,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,1})$. Thus, $i+1$-layer residual-based CNN is:
@@ A.3 The DUAT Format of Residual-based CNNs
where $i=0,1,2,\ldots$. For $i=0$, we have $\mathbf{b}^{\prime}_{1,1}=\mathbf{b}^{\prime}_{1,1}$; and for $i>0$, $\mathbf{b}^{\prime}_{j,1}$ is defined by $\mathbf{b}^{\prime}_{j,1}=\left(\mathbf{W}^{\prime}_{j,1}\mathbf{b}^{\prime}_{j-1,2}+\mathbf{b}^{\prime}_{j,1}\right)+\mathbf{W}^{\prime}_{j,1}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,2}\sigma\left(\mathbf{W}^{\prime}_{k,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,1}\right).$ Thus, $\mathbf{b}^{\prime}_{j,1}$ can be seen as a $j$-layer UAT and the input is $\mathbf{x}_{0}$. For convenience, we also consider the first term as part of the UAT. Additionally, for the term $\mathbf{b}^{\prime}_{j+1,2}$ at layer $i$, it is given by $\mathbf{b}^{\prime}_{j+1,2}=\mathbf{b}^{\prime}_{1,2}\cdots\mathbf{b}^{\prime}_{j,2}+\mathbf{b}^{\prime}_{j+1,2},j=1,2\cdots i$. Naturally, in a multilayer network, $\mathbf{b}^{\prime}_{i,2},i\geq 1$ within the corresponding UAT will also change. However, as shown in the derivation above, these adjustments involve only matrix addition of predefined parameters. Therefore, they can essentially be understood as fixed parameters and do not fundamentally alter the mathematical form of the UAT. Consequently, we will not elaborate on the specific forms of these parameters here. Some examples are provided in Figure 11. Since $\mathbf{b}^{\prime}_{j,1}$ adapts dynamically to the input, the multilayer convolutional network based on residuals can essentially be understood as a DUAT function.
@@
Figure 11: Some examples of the UAT format of multi-layer residual-based convolution. The changes of parameters are represented within the dashed boxes. The parameters on the right of the equations indicate the original values, while those on the left represent the transformed values. There is no specific order of calculation for the parameters within each dashed box, but there is a top-to-bottom calculation order between different dashed boxes.
A.4 The DUAT Format of Transformer-based ViTs
Similarly, to prove that Transformer-based ViTs are also DUAT functions, we start by deriving their general form based on Figure 9 in the main text. Transformers involve two key operations: Multi-Head Attention (MHA) and the Feed-Forward Network (FFN). In matrix-vector form, these operations correspond to Eqs. 19 and 20.
$MHA(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (19)
$FFN(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (20)
@@ A.4 The DUAT Format of Transformer-based ViTs
Thus, a general term in a Transformer-based network can be expressed as Eq. 21. Letting $\mathbf{W}^{\prime}_{i+1,2}=\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}$, we can rewrite the general term as follows:
Clearly, Eq. 22 matches the mathematical form presented in Section A.2: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$. Therefore, a multilayer Transformer-based ViT is indeed DUAT function.
To clearly show the DUAT format of a multilayer ViT, we use the approach outlined in Section A.2. Assuming that the overall mathematical form of $\mathbf{x}^{\prime}_{i}$ aligns with the UAT structure, we decompose $\mathbf{x}^{\prime}_{i}$ into a main term and a residual term, expressed as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,3})+UAT^{R}_{i}$, where $UAT_{i}^{R}=\sum_{j=1}^{i}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ and $i=1,2,\ldots$. Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we get Eq. 23.
@@
Let $\mathbf{W}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}$, $\mathbf{b}^{\prime}_{i+1,3}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})$, $UAT^{R}_{i}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$, $\mathbf{W}^{\prime}_{i+1,2}=\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}$, and $\mathbf{b}^{\prime}_{i+1,2}=(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}$. Substituting these, we obtain $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,3})+UAT^{R}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,2})$. Since $UAT^{R}_{i}$ overall aligns with the UAT, the overall mathematical form of the $i+1$ layers’ ViT also conforms to the UAT.
So the $i+1$ layers’ ViT can be expressed as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (24)
When $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$. We define the parameters for cases where $i\geq 1$. For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$.
Additionally, for $j=1,2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (25)
$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
It means that the whole computing process is serial, if we want to compute $\mathbf{W}^{\prime}_{3,1}$, we must calculate $\mathbf{W}^{\prime}_{2,1}$, because $\mathbf{W}^{\prime}_{3,1}=\mathbf{W}^{\prime}_{3,1}\mathbf{W}^{\prime}_{2,1}$. Thus, we have established that a multilayer ViT is also a DUAT function. The mathematical representation of the ViT aligns with the DUAT framework, with $\mathbf{b}^{\prime}_{j,2}$ closely approximated by DUAT. The most notable difference between ViTs and residual-based CNNs is the dynamic input-dependent nature of the parameters in the MHA mechanism. Consequently, in the DUAT formulation associated with ViT, the parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input.
Figure 12: Some examples of the UAT format of multi-layer Transformer. The changes of parameters are represented within the dashed boxes. The parameters on the right of the equations indicate the original values, while those on the left represent the transformed values. There is no specific order of calculation for the parameters within each dashed box, but there is a top-to-bottom calculation order between different dashed boxes.